.NEED ANSWER ASAP / ANSWER NEVER USED BEFORE

a.)

Maturity Risk Premium

The real risk-free rate is 3%, and inflation is expected to be 3% for the next 2 years. A 2-year Treasury security yields 7.3%. What is the maturity risk premium for the 2-year security? Round your answer to one decimal place.

  %

b.)

Constant Dividend Growth Valuation

Boehm Incorporated is expected to pay a $1.60 per share dividend at the end of this year (i.e., D1 = $1.60). The dividend is expected to grow at a constant rate of 3% a year. The required rate of return on the stock, rs, is 10%. What is the estimated value per share of Boehm's stock? Do not round intermediate calculations. Round your answer to the nearest cent.

c.)

Nonconstant Dividend Growth Valuation

A company currently pays a dividend of $3.2 per share (D0 = $3.2). It is estimated that the company's dividend will grow at a rate of 15% per year for the next 2 years, and then at a constant rate of 7% thereafter. The company's stock has a beta of 1.3, the risk-free rate is 10%, and the market risk premium is 5.5%. What is your estimate of the stock's current price? Do not round intermediate calculations. Round your answer to the nearest cent.

$  

d.)

Return on Common Stock

You buy a share of The Ludwig Corporation stock for $22.10. You expect it to pay dividends of $1.02, $1.0781, and $1.1396 in Years 1, 2, and 3, respectively, and you expect to sell it at a price of $26.10 at the end of 3 years.

  1. Calculate the growth rate in dividends. Round your answer to two decimal places.

      %

  2. Calculate the expected dividend yield. Round your answer to two decimal places.

      %

  3. Assuming that the calculated growth rate is expected to continue, you can add the dividend yield to the expected growth rate to obtain the expected total rate of return. What is this stock's expected total rate of return (assume market is in equilibrium with the required rate of return equal to the expected return)? Do not round intermediate calculations. Round your answer to two decimal places.

      %
Answer #1

a) Maturity Risk Premium- the formula for calculating the maturity risk premium is-Treasurysecurity yield = Realrisk free+Intrate +marketrisk premium

7.3%=3%+3%+x

therefore Market risk Premium(x)=7.3%-6%=1.3%

b)Valuepershare = D1/(rs -9)

i.e. =$1.6/(0.1-0.03)=$22.86

c) To calculate the stock price, firstly we need to calculate the require rate of return for the stock

requiredr ateofreturn(re) = rf + beta * (risk premium)

i.e required rate of return=0.1+1.3*(0.55)=17%

Now, value of the stock=Dividend in the next period/required rate of return. Therefor the value of the given stock is

D0 $       3.20
D1 3.68
D2 4.232
D3 4.52824
Value of the stock(D3/re) 26.40373

d)(1)

Dividend growth
d1 $                                1.02
d2 $                                1.08 0.056961
d3 $                                1.14 0.057045
Dividend growth rate 6% (avg of above)

d)(2) Dividend yield = Share price/ Total dividend =$26.1/$3.24= 8.06

d)(3) There for the expected rate of return= Dividend growth rate+Dividend yield=6+8.06=14.06%